Two parallel straight long conducting wires which are placed at a distance R from each other are carrying equal currents I in opposite directions. The value of magnetic induction at a point situated at a point situated X from one wire in between the wires will be..?

Grade:10

1 Answers

Eshan
5 years ago
Dear student,

Magnetic field at the point due to one of the wires=B_1=\dfrac{\mu_0I}{2\pi x}

Distance of second wire from the point=R-x
Hence magnetic field at the same point due to the other wire=B_2=\dfrac{\mu_0 i}{2\pi (R-x)}
The direction of this magnetic field wil be same as that of first wire.
HenceB_T=B_1+B_2=\dfrac{\mu_0i}{2\pi}(\dfrac{1}{x}+\dfrac{1}{R-x})